38. Determine the structure based on the following molecular formula, IR, and NMR spectrum: Formula, Coloạo 2H2H
Answer #1

Go AIO DBE = 10+1 -14 = 11-7-4 on given data proposed Structure Dropoled Bused CH3 om -OH the СЛt3 CA IHNMR ICH3 CH3 X CH3 at